OpenStudy (anonymous):

sqrt-54+sqrt-96 plz help im stuck on this

OpenStudy (anonymous):

i have to simplify the given expression

OpenStudy (anonymous):

54/2=27 27/3=9 9/3=3 3/3=1 so \[sqrt(54)=sqrt(2*3^2*3)\] We can take the square root of \[3^2=3\] \[sqrt(54)=3*sqrt(3*2)=3*sqrt(6)\] Try thesame for the other one.

OpenStudy (anonymous):

-96/2=-48 -48/2=-24 -24/2=-12 -12/2=-6 -6/2=-3 -3 is my answer is that correct

OpenStudy (anonymous):

Yes, so now you know that \[sqrt(96)=sqrt(2^5*3)\] We might not be able to take the square root of 2^5, but we can of 2^4. So we get \[sqrt(96)=sqrt(2^4*2*3)\] I'm sure you can do the rest :)

OpenStudy (anonymous):

Wait

OpenStudy (anonymous):

sqrt 27 + sqrt 24 3 sqrt of 3 + 2 sqrt of 6 2. sqrt 216t + sqrt 96t (sqrt 36 * sqrt 6t) + (sqrt 16 * sqrt 6t) 6 sqrt of 6t + 4 sqrt of 6t 10 sqrt of 6t

OpenStudy (anonymous):

i can do the rest thanks for your help